TERAHERTZ VIBRATION-ROTATION-TUNNELING (VRT) SPECTROSCOPY OF THE WATER TRIMER (D_{2}O)_{3}: COMPLETE CHARACTERIZATION OF THE 2.94 THZ TORSIONAL BAND (k^{n} = ±2^{1} ← 0^{0})

J. -X. HAN, L. K. TAKAHASHI, W. LIN, E. LEE, F. N. KEUTSCH, AND R. J. SAYKALLY, Department of Chemistry, University of California, Berkeley, California 94720.

We report the complete measurement and analysis of the perpendicular k^{n} = ±2^{1} ← 0^{0} (D_{2}O)_{3} torsional band with an origin of 2940.9376(3) GHz. The upper state is the highest torsional state yet observed. All previously known torsional transitions were included in a new global analysis of the six observed torsional bands, using the effective Hamiltonians derived by van der Avoid et al.